# Position: The Hidden Costs and Measurement Gaps of Reinforcement Learning with Verifiable Rewards

### v1

@@ 7 Conclusion
RLVR delivers real gains on verifiable tasks, but the field often over-indexes on headline accuracy while under-weighting taxes (hallucination/overconfidence, instruction drift, safety & privacy exposure) and measurement (budget mismatch, judge fragility, calibration drift, contamination). Under parity controls and calibration tracking, several celebrated “reasoning gains” shrink, suggesting that part of the progress reflects distributional sharpening rather than durable expansion.
@@
Broader relevance. The same pressures arise in SFT, RLHF, and test-time compute (e.g., CoT and maj@k). Our protocol—budget parity with saturation curves, calibration/abstention metrics, judge robustness checks, and contamination audits—applies unchanged beyond RLVR.
@@
Why focus on RLVR. We center RLVR because programmatic verifiers make the measurement problem tractable and reveal where gains come from. As a result, it is a useful proving ground for reliability methods that transfer to broader LLM training and deployment. We hope this synthesis helps the community separate genuine capability expansion from artifacts of budgets, metrics, and data, and encourages tax-aware training and reporting across RLVR, SFT, and RLHF.
Ethics Statement
This paper raises some potential ethics concerns under the ICLR Code of Ethics and describes the steps we took to mitigate them:
Privacy & leakage. Long, explicit chains-of-thought can reveal sensitive content or training artifacts. We evaluate only on public math/code benchmarks and do not use private prompts. For any illustrative traces, we sanitize identifiers and redact potentially sensitive strings. Our standardized evaluation caps best-of-$N$ and CoT length to reduce leakage surfaces.
Dual-use: jailbreaks & misuse. Methodological details about jailbreak stress tests could be repurposed to increase harm. We report aggregate metrics and robustness deltas, but do not release attack payloads or automation scripts that would materially increase misuse without corresponding defenses. Our evaluation protocol co-optimizes calibrated abstention alongside accuracy.
Measurement integrity. Selective budgeting, fragile judge pipelines, or metric switching can misrepresent results. To reduce this risk, we use matched decoding budgets, fixed verifier/prompt families, multi-seed averaging, and release configurations (decoding parameters, seeds, templates) sufficient to reproduce our reported numbers.
Reproducibility Statement
We make our results reproducible by fixing verifiers and a shared prompt family (section 4), matching decoding budgets and reporting decoding parameters, and averaging over three seeds (Appendix 4, Table2; Figure 2). Datasets and methodology for contamination screens and judge-robustness probes referenced in section 5 are provided with instructions to re-run the audits. Hardware details and aggregate compute ($128{\times}$H100 (96 GB) cluster and $\approx$3,500 GPU-hours) are included to support cost and throughput replication.
@@ Appendix A Appendix
pass@k = probability at least one of $k$ samples is correct; avg@k = mean pass@1 across $k$ draws; maj@k = majority vote over $k$; ECE = expected calibration error. Arrows ($\uparrow$/↓) indicate whether higher/lower is better.
Usage of Large Language Models.
We used an LLM assistant as a productivity tool for light editing and formatting: smoothing wording the authors had written, polishing captions, refactoring tables/macros, and drafting small utility snippets (e.g., plotting and CSV loaders) that we then modified and verified. All research ideas, evaluation designs, hyperparameters, and analysis decisions are by the authors, and every experiment and number reported was run and checked by us.
Appendix B Potential Criticisms and Responses
@@
RLVR can expand capability under specific regimes (e.g., prolonged RL with resets and KL control; unlikeliness-based shaping; domain-targeted RLVR with sparse pretraining) (Liu et al., 2025a; He et al., 2025a; Cheng et al., 2025b). Our claim is not negation but measurement: without tax-aware training and parity-controlled evaluation, expansion is often overstated. The protocol in §6 is compatible with these positive regimes while curbing overclaiming.
@@
Data quality and curricula matter (Chen et al., 2025a; Wen et al., 2025a), but reward/metric design independently shapes failure modes (determinacy/overconfidence under judge rewards; terse/off-topic outputs under naive factuality; calibration drift) (Chen et al., 2025c; Leng et al., 2024; Hochlehnert et al., 2025). We treat both data and objectives as first-class knobs (see §6).
@@
We use pass@k for practical relevance, but ask for matched budgets and saturation curves. Under these controls, headline gaps often shrink or flip (Hochlehnert et al., 2025). Reporting AUC alongside pass@k keeps comparisons fair (§4).
@@
Jailbreak success scales with attempts/context; many-shot tuning can reduce measured rates in controlled settings (Zhou et al., 2025; Jiang et al., 2025; Ackerman & Panickssery, 2025). Longer CoT increases leakage surfaces (Green et al., 2025). Our protocol caps Best-of-NN/CoT length at eval and co-optimizes abstention/privacy with accuracy (§6).
@@
When verifiers are unavailable or partial, we use LLM judges with robustness probes (prompt/format perturbations, adversarial tests) and publish templates/configs; judges are manipulable, so robustness deltas are reported (Zhao et al., 2025b; Sun et al., 2025a). Prefer verifiers when possible (§4).
@@
Appendix C Standardized Evaluations
We report full results under matched budgets, fixed verifiers, and a shared prompt family. These tables extend Section 4 and underlie the saturation-curve argument: once $k$, templates, and dataset versions are controlled, several headline gaps narrow. All results in this section use the same verifier and prompt template family; we specify $k$, number of seeds, and dataset versions below.
@@ Appendix C Standardized Evaluations
Appendix D Sequence Length and Compute Footprint
We summarize generated token lengths by benchmark as a proxy for compute footprint and exposure surface. Longer chains increase latency and may worsen calibration and safety/privacy risks (Section 3), So we report typical generation lengths under the same decoding budgets used in Appx. C.
Table 5: Median generated output length (tokens per problem; answer + reasoning) under the standardized decoding setup. Context: Compute/privacy footprint complementing accuracy. Setup: Same prompts/verifier as Appx. C; decoding budget k=32k{=}32; lengths aggregated across all samples per problem (correct and incorrect). Takeaway: Reasoning-tuned models often produce much longer traces (e.g., 2–8× vs. instruct baselines), which impacts evaluation cost and safety/privacy exposure.
@@ Appendix D Sequence Length and Compute Footprint
DeepSeek-R1-Distill-Qwen-32B 10979 13012 5826 3652 4663 7924 7676
@@
Appendix E Data Contamination Probes: Math
@@
We use partial-prompt completion to detect tail reconstruction: reveal the first $x\%$ of the problem ($x\!\in\!\{80,60,40\}$), greedily decode the remainder, and score ACC/EM/ROUGE-L at that prefix. High scores at large prefixes indicate memorization of legacy sets rather than reasoning. Tables here expand Section 5.
@@
Table 6: Accuracy of Qwen3 checkpoints on five mathematics benchmarks when each model receives only the first $x\%$ of the question (x=80,60,40x\!=\!80,60,40) and must greedily complete the remainder. Columns “ACC (X%)” report the average accuracy at that prefix length. Takeaway: Qwen3 variants achieve high ACC@80 on legacy MATH-500/AMC-23 but collapse on AIME-2025, consistent with contamination in older sets.
Model Dataset ACC (80%) ACC (60%) ACC (40%)
MATH‑500 26.40 14.40 5.40
AMC‑23 20.00 5.00 0.00
Qwen3‑0.6B‑Base AIME 2024 3.33 0.00 0.00
AIME 2025 0.00 0.00 0.00
Minerva‑Math 1.10 0.00 0.00
MATH‑500 33.80 18.40 9.00
AMC‑23 20.00 15.00 0.00
Qwen3‑1.7B‑Base AIME 2024 0.00 0.00 0.00
AIME 2025 3.33 3.33 0.00
Minerva‑Math 2.94 2.21 0.74
MATH‑500 43.20 28.20 15.00
AMC‑23 37.50 37.50 17.50
Qwen3-4B‑Base AIME 2024 10.00 6.67 10.00
AIME 2025 3.33 0.00 0.00
Minerva‑Math 4.78 2.31 0.74
MATH‑500 52.00 35.60 24.80
AMC‑23 35.00 25.00 27.50
Qwen3‑8B‑Base AIME 2024 13.33 6.67 10.00
AIME 2025 3.33 3.33 0.00
Minerva‑Math 5.88 3.31 1.84
MATH‑500 58.20 44.80 30.20
AMC‑23 47.50 37.50 32.50
Qwen3‑14B‑Base AIME 2024 10.00 16.67 16.67
AIME 2025 0.00 0.00 0.00
Minerva‑Math 5.88 2.21 2.94
MATH‑500 62.40 51.40 37.80
AMC‑23 40.00 25.00 25.00
Qwen3‑30B‑A3B AIME 2024 13.33 23.33 13.33
AIME 2025 3.33 0.00 0.00
Minerva‑Math 4.78 3.21 1.84
@@
Table 7: Accuracy of Qwen2.5 and Llama‑3 checkpoints on five mathematics benchmarks when each model receives only the first $x\%$ of the question (x=80,60,40x\!=\!80,60,40) and must greedily complete the remainder. Takeaway: Qwen2.5 shows strong tail reconstruction on legacy math; Llama-3-1.8B remains near zero across sets, reinforcing the contamination interpretation.
Model Dataset ACC (80%) ACC (60%) ACC (40%)
MATH‑500 58.00 44.40 29.20
AMC‑23 52.50 42.50 32.50
Qwen2.5‑Math‑7B AIME 2024 16.67 20.00 16.67
AIME 2025 0.00 0.00 0.00
Minerva‑Math 7.72 4.41 2.94
MATH‑500 44.80 27.80 15.00
AMC‑23 27.50 27.50 22.50
Qwen2.5‑7B AIME 2024 6.67 0.00 3.33
AIME 2025 3.33 3.33 0.00
Minerva‑Math 6.62 4.78 1.47
MATH‑500 44.80 28.60 14.00
AMC‑23 40.00 15.00 12.50
Qwen2.5‑7B‑Instruct AIME 2024 3.33 0.00 0.00
AIME 2025 3.33 0.00 0.00
Minerva‑Math 10.29 5.51 3.31
MATH‑500 50.60 35.80 21.00
AMC‑23 40.00 27.50 27.50
Qwen2.5‑14B AIME 2024 10.00 3.33 6.67
AIME 2025 3.33 0.00 0.00
Minerva‑Math 8.46 6.25 2.21
MATH‑500 60.00 47.80 32.00
AMC‑23 52.50 45.00 42.50
Qwen2.5‑32B AIME 2024 16.67 13.33 10.00
AIME 2025 3.33 0.00 6.67
Minerva‑Math 9.93 4.41 2.94
MATH‑500 2.80 2.40 2.00
AMC‑23 0.00 0.00 0.00
Llama‑3‑1.8B AIME 2024 0.00 0.00 0.00
AIME 2025 0.00 0.00 0.00
Minerva‑Math 1.84 0.37 0.00
@@
Table 8: RougeL and exact‑match (EM) scores for Qwen3.0 checkpoints on five mathematics benchmarks when each model receives only the first $x\%$ of the question (x=80,60,40x\!=\!80,60,40) and must greedily complete the remainder. Columns “R@xx” and “EM@xx” report the average ROUGE‑L and EM for that prefix length. Takeaway: Elevated ROUGE-L/EM at large prefixes on legacy sets (but not on AIME-2025) indicates suffix reconstruction rather than emergent reasoning.
@@
Table 9: RougeL and exact‑match (EM) scores for Qwen2.5 and Llama‑3 on five mathematics benchmarks when each model receives only the first $x\%$ of the question (x=80,60,40x\!=\!80,60,40) and must greedily complete the remainder. Takeaway: Qwen2.5 shows strong reconstruction on legacy math; Llama-3 remains low.
@@
MATH‑500 79.63 61.60 70.30 41.20 60.81 25.40
AMC‑23 77.35 57.50 71.47 50.00 66.29 42.50
Qwen2.5‑Math‑7B AIME 2024 69.35 53.33 55.90 30.00 56.53 30.00
@@
Minerva‑Math 33.42 2.94 30.41 0.37 26.88 0.00
MATH‑500 66.09 38.20 58.15 18.00 48.34 6.00
AMC‑23 69.47 47.50 67.70 45.00 61.08 35.00
Qwen2.5‑7B AIME 2024 61.28 30.00 46.19 16.67 46.14 16.67
@@
Minerva‑Math 33.88 2.94 31.13 0.00 27.06 0.00
MATH‑500 60.05 28.60 51.10 10.40 43.41 2.60
AMC‑23 61.38 30.00 51.22 17.50 42.90 10.00
Qwen2.5‑7B‑Instr AIME 2024 55.41 20.00 36.28 0.00 30.64 0.00
@@
Minerva‑Math 34.37 2.94 28.50 0.00 25.99 0.00
MATH‑500 69.50 45.00 61.20 27.00 51.24 12.80
AMC‑23 74.58 47.50 68.20 42.50 67.82 37.50
Qwen2.5‑14B AIME 2024 65.42 40.00 49.76 23.33 48.97 23.33
@@
Minerva‑Math 35.64 3.68 31.30 0.00 27.93 0.00
MATH‑500 76.98 56.60 71.14 41.20 58.32 25.20
AMC‑23 79.18 62.50 74.59 60.00 75.78 55.00
Qwen2.5‑32B AIME 2024 73.25 56.67 60.92 40.00 60.47 33.33
@@
Minerva‑Math 34.88 2.94 32.92 0.00 28.32 0.00
MATH‑500 46.49 14.00 38.75 3.00 32.19 0.60
AMC‑23 40.46 0.00 29.87 0.00 25.28 0.00
Llama‑3‑1.8B AIME 2024 49.70 16.67 31.45 0.00 27.13 0.00
@@
Minerva‑Math 35.32 1.47 29.17 0.00 26.95 0.00
Appendix F Data Contamination Probes: SimpleQA (Control)
To test whether partial-prompt reconstruction reflects a general suffix-completion ability, we repeat the probe on SimpleQA. If math-set effects were general, we would expect similarly high ROUGE-L/EM at large prefixes here. They are not: scores cluster modestly across families.
Table 10: Rouge‑L and exact‑match (EM) scores on SimpleQA when each model receives only the first $x\%$ of the question (x=80,60,40x\!=\!80,60,40) and must greedily complete the remainder. Takeaway: No consistent Qwen advantage; effects seen on legacy math do not generalize, supporting the contamination interpretation.

### v2

@@ 7 Conclusion
Our central position is that RLVR is effective and deployable for verifiable domains, but headline “reasoning” gains should be treated as provisional unless a small core of tax-aware controls (budget parity, calibration/abstention tracking, robust evaluation with at least one judge stress test, and a simple contamination audit) is enforced. RLVR delivers real gains on verifiable tasks, but the field often over-indexes on headline accuracy while under-weighting taxes and measurement. Under budget parity controls and calibration tracking, several celebrated “reasoning gains” shrink, suggesting that part of the progress reflects distributional sharpening rather than durable expansion.
@@
Broader relevance. The same pressures arise in SFT, RLHF, and test-time compute (e.g., CoT and maj@k). This tax-aware minimum standard applies unchanged beyond RLVR.
@@
Why focus on RLVR. We focus on RLVR because verifiers make these trade-offs measurable and make RLVR a useful proving ground. As a result, it is a useful proving ground for reliability methods that transfer to broader LLM training and deployment. We hope this synthesis helps the community separate genuine capability expansion from artifacts of budgets, metrics, and data, and encourages tax-aware training and reporting across RLVR, SFT, and RLHF.
Limitations
Our goal is measurement clarity rather than a complete theory of when RLVR yields capability expansion, so our evidence has clear boundaries. First, our empirical analyses focus on verifiable domains (math, code, structured QA) using open checkpoints and public benchmarks; findings may not transfer directly to agentic, multimodal, or tool-use settings where programmatic verification is harder. Second, our attempt inflation and calibration analyses depend on operationalizing "attempts" (e.g., extractable answers versus explicit refusals); different extraction rules shift absolute rates, so we emphasize shared-item comparisons to reduce sensitivity. Third, our partial-prompt contamination probe is a high-precision indicator of tail reconstruction on some datasets but is not exhaustive: it can miss semantic overlap and does not quantify training-set membership. We treat it as one triangulation signal alongside provenance checks and fresh held-out sets. We omit approximate nearest-neighbor or fuzzy-hash matching against training corpora and shard-level provenance analysis, which would strengthen contamination claims but require additional infrastructure and training-data access. Finally, we exclude RLVR reward-component ablations (e.g., explicit refusal, grounding, or calibration terms) because they require additional training runs; we view them as promising next steps.
Ethical Considerations
This paper is primarily about evaluation and reporting practices, but it touches on safety and privacy. Longer reasoning traces can increase the chance of exposing sensitive strings or unsafe intermediate content. To mitigate risk, we restrict experiments to public benchmarks and report aggregate statistics rather than releasing jailbreak payloads or automated attack scripts. We also highlight calibration, abstention, and provenance as first-class reporting targets because inflated headline accuracy can create misleading impressions of real-world reliability. Finally, our recommendations can increase evaluation cost (e.g., multi-seed, budget-matched saturation curves); we view this as a necessary trade-off for measurement integrity when results are used to justify deployment.
Acknowledgments
This work was supported in part by Google Cloud credits through the Gemini Academic Program, which enabled large-scale evaluation and experimentation. We also thank DSO National Laboratories (Singapore) for support on knowledge-focused reasoning research. The views and conclusions expressed in this paper are those of the authors and do not necessarily reflect the official policies of the sponsors.
@@ Appendix A Appendix
pass@k = probability at least one of $k$ samples is correct; avg@k = mean pass@1 across $k$ draws; maj@k = majority vote over $k$; ECE = expected calibration error. Arrows ($\uparrow$$\downarrow$) indicate whether higher or lower is better.
A.1 Supplementary Figure for the Introduction
This appendix subsection contains Figure 2, which shows the RLVR activity trend alongside AIME performance.
Figure 2: Monthly RLVR activity vs. AIME performance (time span: May 2024–June 2025). Left axis: count of pages per month whose title or abstract contains “RLVR” or “reinforcement learning with verifiable rewards” (Google Scholar and arXiv via SerpAPI). Right axis: pass@1 (%) on AIME-24 (circles) and AIME-25 (squares) for selected models; labels show model names.
A.2 Potential Criticisms and Responses
@@
RLVR can expand capability under specific regimes (e.g., prolonged RL with resets and KL control; unlikeliness-based shaping; domain-targeted RLVR with sparse pretraining) (Liu et al., 2025a; He et al., 2025a; Cheng et al., 2025). Our claim is not negation but measurement: without tax-aware training and parity-controlled evaluation, expansion is often overstated. The protocol in section 6 is compatible with these positive regimes while curbing overclaiming.
@@
Data quality and curricula matter (Chen et al., 2025a; Wen et al., 2025a), but reward/metric design independently shapes failure modes (determinacy/overconfidence under judge rewards; terse/off-topic outputs under naive factuality; calibration drift) (Chen et al., 2025b; Leng et al., 2024; Hochlehnert et al., 2025). We treat both data and objectives as first-class knobs (see section 6).
@@
We use pass@k for practical relevance, but ask for matched budgets and saturation curves. Under these controls, headline gaps often shrink or flip (Hochlehnert et al., 2025). Reporting AUC alongside pass@k keeps comparisons fair (section 4).
@@
Jailbreak success scales with attempts/context; many-shot tuning can reduce measured rates in controlled settings (Zhou et al., 2025a; Jiang et al., 2025; Ackerman and Panickssery, 2025). Longer CoT increases leakage surfaces (Green et al., 2025; Zhou et al., 2025b). Our protocol caps Best-of-NN/CoT length at eval and co-optimizes abstention/privacy with accuracy (section 6).
@@
When verifiers are unavailable or partial, we use LLM judges with robustness probes (prompt/format perturbations, adversarial tests) and publish templates/configs; judges are manipulable, so robustness deltas are reported (Zhao et al., 2025; Sun et al., 2025a). Prefer verifiers when possible (section 4).
@@
A.3 Standardized Evaluations
We report full results under matched budgets, fixed verifiers, and a shared prompt family. These tables extend section 4 and underlie the saturation-curve argument: once $k$, templates, and dataset versions are controlled, several headline gaps narrow. See Table 4 below for the full standardized benchmark results. All results in this section use the same verifier and prompt template family; we specify $k$, number of seeds, and dataset versions below.
@@ A.3 Standardized Evaluations
A.4 Sequence Length and Compute Footprint
We summarize generated token lengths by benchmark as a proxy for compute footprint and exposure surface. Longer chains increase latency and may worsen calibration and safety/privacy risks (section 3), so we report typical generation lengths under the same decoding budgets used in Appendix A.3. See Table 5 below for the token-length breakdown by model and benchmark.
[…]
@@ A.4 Sequence Length and Compute Footprint
DeepSeek-R1-Distill-Qwen-32B 10979 13012 5826 3652 4663 7924 $76$
@@
A.5 Data Contamination Probes: Math
@@
We use partial-prompt completion to detect tail reconstruction: reveal the first $x\%$ of the problem ($x\in\{80,60,40\}$), greedily decode the remainder, and score ACC/EM/ROUGE-L at that prefix. High scores at large prefixes indicate memorization of legacy sets rather than reasoning. Tables here expand section 5. Accuracy results are reported in Tables 6 and 7; text-overlap and exact-match proxies are reported in Tables 8 and 9.
@@
Table 6: Accuracy of Qwen3 checkpoints on five mathematics benchmarks when each model receives only the first $x\%$ of the question ($x=80,60,40$) and must greedily complete the remainder. Columns “ACC ($x\%$)” report the average accuracy at that prefix length. Takeaway: Qwen3 variants achieve high ACC@80 on legacy MATH-500/AMC-23 but collapse on AIME-2025, consistent with contamination in older sets.
Model Dataset ACC (80%) ACC (60%) ACC (40%) MATH-500 26.40 14.40 5.40 AMC-23 20.00 5.00 0.00 Qwen3-0.6B-Base AIME 2024 3.33 0.00 0.00 AIME 2025 0.00 0.00 0.00 Minerva-Math 1.10 0.00 0.00 MATH-500 33.80 18.40 9.00 AMC-23 20.00 15.00 0.00 Qwen3-1.7B-Base AIME 2024 0.00 0.00 0.00 AIME 2025 3.33 3.33 0.00 Minerva-Math 2.94 2.21 0.74 MATH-500 43.20 28.20 15.00 AMC-23 37.50 37.50 17.50 Qwen3-4B-Base AIME 2024 10.00 6.67 10.00 AIME 2025 3.33 0.00 0.00 Minerva-Math 4.78 2.31 0.74 MATH-500 52.00 35.60 24.80 AMC-23 35.00 25.00 27.50 Qwen3-8B-Base AIME 2024 13.33 6.67 10.00 AIME 2025 3.33 3.33 0.00 Minerva-Math 5.88 3.31 1.84 MATH-500 58.20 44.80 30.20 AMC-23 47.50 37.50 32.50 Qwen3-14B-Base AIME 2024 10.00 16.67 16.67 AIME 2025 0.00 0.00 0.00 Minerva-Math 5.88 2.21 2.94 MATH-500 62.40 51.40 37.80 AMC-23 40.00 25.00 25.00 Qwen3-30B-A3B AIME 2024 13.33 23.33 13.33 AIME 2025 3.33 0.00 0.00 Minerva-Math 4.78 3.21 1.84
@@
Table 7: Accuracy of Qwen2.5 and Llama-3 checkpoints on five mathematics benchmarks when each model receives only the first $x\%$ of the question ($x=80,60,40$) and must greedily complete the remainder. Takeaway: Qwen2.5 shows strong tail reconstruction on legacy math; Llama-3-1.8B remains near zero across sets, reinforcing the contamination interpretation.
Model Dataset ACC (80%) ACC (60%) ACC (40%) MATH-500 58.00 44.40 29.20 AMC-23 52.50 42.50 32.50 Qwen2.5-Math-7B AIME 2024 16.67 20.00 16.67 AIME 2025 0.00 0.00 0.00 Minerva-Math 7.72 4.41 2.94 MATH-500 44.80 27.80 15.00 AMC-23 27.50 27.50 22.50 Qwen2.5-7B AIME 2024 6.67 0.00 3.33 AIME 2025 3.33 3.33 0.00 Minerva-Math 6.62 4.78 1.47 MATH-500 44.80 28.60 14.00 AMC-23 40.00 15.00 12.50 Qwen2.5-7B-Instruct AIME 2024 3.33 0.00 0.00 AIME 2025 3.33 0.00 0.00 Minerva-Math 10.29 5.51 3.31 MATH-500 50.60 35.80 21.00 AMC-23 40.00 27.50 27.50 Qwen2.5-14B AIME 2024 10.00 3.33 6.67 AIME 2025 3.33 0.00 0.00 Minerva-Math 8.46 6.25 2.21 MATH-500 60.00 47.80 32.00 AMC-23 52.50 45.00 42.50 Qwen2.5-32B AIME 2024 16.67 13.33 10.00 AIME 2025 3.33 0.00 6.67 Minerva-Math 9.93 4.41 2.94 MATH-500 2.80 2.40 2.00 AMC-23 0.00 0.00 0.00 Llama-3-1.8B AIME 2024 0.00 0.00 0.00 AIME 2025 0.00 0.00 0.00 Minerva-Math 1.84 0.37 0.00
@@
Table 8: RougeL and exact-match (EM) scores for Qwen3.0 checkpoints on five mathematics benchmarks when each model receives only the first $x\%$ of the question ($x=80,60,40$) and must greedily complete the remainder. Columns “R@x” and “EM@x” report the average ROUGE-L and EM for that prefix length. Takeaway: Elevated ROUGE-L/EM at large prefixes on legacy sets (but not on AIME-2025) indicates suffix reconstruction rather than emergent reasoning.
@@
Table 9: RougeL and exact-match (EM) scores for Qwen2.5 and Llama-3 on five mathematics benchmarks when each model receives only the first $x\%$ of the question ($x=80,60,40$) and must greedily complete the remainder. Takeaway: Qwen2.5 shows strong reconstruction on legacy math; Llama-3 remains low.
@@
MATH-500 79.63 61.60 70.30 41.20 60.81 25.40
AMC-23 77.35 57.50 71.47 50.00 66.29 42.50
Qwen2.5-Math-7B AIME 2024 69.35 53.33 55.90 30.00 56.53 30.00
@@
Minerva-Math 33.42 2.94 30.41 0.37 26.88 0.00
MATH-500 66.09 38.20 58.15 18.00 48.34 6.00
AMC-23 69.47 47.50 67.70 45.00 61.08 35.00
Qwen2.5-7B AIME 2024 61.28 30.00 46.19 16.67 46.14 16.67
@@
Minerva-Math 33.88 2.94 31.13 0.00 27.06 0.00
MATH-500 60.05 28.60 51.10 10.40 43.41 2.60
AMC-23 61.38 30.00 51.22 17.50 42.90 10.00
Qwen2.5-7B-Instr AIME 2024 55.41 20.00 36.28 0.00 30.64 0.00
@@
Minerva-Math 34.37 2.94 28.50 0.00 25.99 0.00
MATH-500 69.50 45.00 61.20 27.00 51.24 12.80
AMC-23 74.58 47.50 68.20 42.50 67.82 37.50
Qwen2.5-14B AIME 2024 65.42 40.00 49.76 23.33 48.97 23.33
@@
Minerva-Math 35.64 3.68 31.30 0.00 27.93 0.00
MATH-500 76.98 56.60 71.14 41.20 58.32 25.20
AMC-23 79.18 62.50 74.59 60.00 75.78 55.00
Qwen2.5-32B AIME 2024 73.25 56.67 60.92 40.00 60.47 33.33
@@
Minerva-Math 34.88 2.94 32.92 0.00 28.32 0.00
MATH-500 46.49 14.00 38.75 3.00 32.19 0.60
AMC-23 40.46 0.00 29.87 0.00 25.28 0.00
Llama-3-1.8B AIME 2024 49.70 16.67 31.45 0.00 27.13 0.00
@@
Minerva-Math 35.32 1.47 29.17 0.00 26.95 0.00
A.6 Data Contamination Probes: SimpleQA (Control)
To test whether partial-prompt reconstruction reflects a general suffix-completion ability, we repeat the probe on SimpleQA. If math-set effects were general, we would expect similarly high ROUGE-L/EM at large prefixes here. They are not: scores cluster modestly across families. See Table 10 below for the ROUGE-L and exact-match results across prefix lengths.
Table 10: Rouge-L and exact-match (EM) scores on SimpleQA when each model receives only the first $x\%$ of the question ($x=80,60,40$) and must greedily complete the remainder. Takeaway: No consistent Qwen advantage; effects seen on legacy math do not generalize, supporting the contamination interpretation.